Calculate the number of iodine atoms in 3.50g of iodine molecules. Give your answer to 3 significant figures. The Avogadro constant is 6.022*1022 mol-1.


1
Expert's answer
2021-03-16T08:17:52-0400

mass of I2 = 3.50 g

molar mass of I2 = 253.81 g/mol


number of moles of I2 = "\dfrac{3.50g}{253.81g/mol} = 0.0138\ moles"


1 mole of Iodine molecule = 2 moles of Iodine atoms

0.0138 moles of Iodine molecules = 0.0276 moles of Iodine atoms


1 mole of Iodine atom = 6.022 × 10²³ atoms

"\therefore" 0.0276 moles of Iodine atom = 1.662 × 10²² atoms